Solve for x
1/8=4^6x+2

Answer by flame8855(424)

2^-3 = 4^(6x+2)
ln both sides
-3ln2 = (6x+2) ln 4 = 2(6x+2)ln2
-3=2(6x+2) = 12x+4
x = -7/12